OpenStudy (anonymous):

solve (x-5)/(3-x)>=0

OpenStudy (anonymous):

x>=5.

myininaya (myininaya):

the top is 0 when x=5 the bottom is 0 when x=3 ------|---------|-------- 3 5 we need to test the intervals: (-inf,3);(3,5);(5,inf) so if we plug in a number btw -inf and 3 like 0 we get f(0)=(0-5)/(3-0)=-5/3 so this not bigger than 0 now we plug in number btw 3 and 5 like 4 and we get f(4)=(4-5)/(3-4)=-1/(-1)=1 this is bigger than 0 so interval (3,5) is apart of our solution and don't forget we include 5 since it says =0 so actually we have (3,5] last interval is (5,inf) we plug in number like 6 and we get f(6)=(6-5)/(3-6)=1/(-3)=-1/3 not bigger than 0 so the solution is (3,5]